A. The following are the heights of four students in centimeters. Suppose


samples of size 2 are taken from this population of four students.


STUDENTS | HEIGHT(in cm)



Cardo 125



Alyana 120



Joaquin 130



Flora 110



a. Compute the mean of the population


b. Compute the mean of the sampling distribution of the sample means

1
Expert's answer
2022-05-02T14:43:29-0400

We have population values 125, 120, 130, 110, population size N=4 and sample size n=2.

Mean of population "(\mu)" = 

"\dfrac{125+120+130+110}{4}=121.25"


The number of possible samples which can be drawn without replacement is "^{N}C_n=^{4}C_2=6."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 125,120 & 245/2 \\
 \hdashline
 2 & 125, 130 & 255/2 \\
 \hdashline
 3 & 125,110 & 235/2 \\
 \hdashline
 4 & 120,130 & 250/2 \\
 \hdashline
 5 & 120,110 & 230/2 \\
 \hdashline
 6 & 130,110 & 240/2 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) \\ \hline
 230/2 & 1/6 & 230/12\\
 \hdashline
 235/2 & 1/6 & 235/12\\
 \hdashline
 240/2 & 1/6 & 240/12\\
 \hdashline
 245/2 & 1/6 & 245/12 \\
 \hdashline
 250/2 & 1/6 & 250/12\\
 \hdashline
 255/2 & 1/6 & 255/12 \\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=121.25=\mu"